OpenStudy (calleyleeann):

What is the speed of a proton accelerated by a potential difference of 250 MV?

OpenStudy (perl):

A proton (q = e , m = 1.67*10^-27 kg ) moves "downhill" through a potential difference of 250 MegaVolts (250 * 10^6 Volts)

OpenStudy (perl):

ok so far?

OpenStudy (calleyleeann):

I'm really confused about the "accelerated by a potential difference of 250 MV" part... :/ I don't understand what it means

OpenStudy (calleyleeann):

and I'm unsure how I can utilize that information

OpenStudy (perl):

The proton moves "downhill" through a 250*10^6 V potential difference. The loss in PE must equal to the gain in KE. PE= q*V = (1.6 * 10^-19 C) (250 *10^6 V) = 4.0*10^-11 J KE = 1/2 * m * v^2 solve PE = KE

OpenStudy (calleyleeann):

What is q?

OpenStudy (perl):

q is the charge of an electron in coulombs

OpenStudy (calleyleeann):

Oh okay. I understand now. Thank you!!

OpenStudy (perl):

I got 2.18 * 10^8 m/s

OpenStudy (calleyleeann):

Alrighty. I'm working on it... slowly but surely